# Quantifying the Dynamics of Harm Caused by Retracted Research

### v1

@@ Discussion
Our findings indicate that the harm caused by retracted papers and their dissemination is closely linked to entrenched ”norms” in academia. Specifically, the excessive emphasis on recent research while neglecting long-term contributions, combined with the reliance on journal impact factors as primary evaluation metrics, prevents stakeholders from effectively addressing the harm caused by retracted papers. In the absence of punitive measures, addressing these enduring ”norms” within the academic system becomes critically important. Our study connects the harm caused by problematic papers to systemic challenges within the academic recognition framework, offering an additional perspective and direction for addressing this issue.
Our study has certain limitations. The forms of harm are not limited to paper citations; they may also include factors such as the likelihood of publication, funding acquisition, and reputation. However, quantifying aspects such as reputation and funding is challenging, and data availability is limited. Meanwhile, the publication and citation of papers are widely recognized as the ”currency” in the academic community and are closely linked to other factors like reputation and funding acquisition (?). Therefore, analyzing the citation patterns of papers can largely reflect the harm caused by retracted papers. Additionally, due to the lack of data on the paper submission process, we were unable to directly connect publication with the harm caused by retracted papers or explore the potential harm associated with the publication process. However, citation count remains crucial in academic evaluation, effectively revealing both the harm caused by retracted papers and the propagation of that harm. Our findings have broad policy implications, offering valuable insights for policy development. Citation-based evaluation mechanisms are currently widespread across various stakeholders (e.g., institutions, authors, journals), influencing reward programs, research funding allocation, award decisions, and even salary and bonus structures, as well as journal rankings (?).
@@
4.
C. Piller, Blots on a field? Science377 (6604), 358–363 (2022).
5.
J. Brainard, J. You, et al., What a massive database of retracted papers reveals about science publishing’s ‘death penalty’. Science25 (1), 1–5 (2018).
6.
H. Else, Biomedical paper retractions have quadrupled in 20 years—why? Nature630 (8016), 280–281 (2024).
10.
M. MAURYA, R. MUNSHI, Reasons for retraction of clinical research articles in PubMed indexed medical journals from 2012 to 2022. Med Ethics9 (4), 296–300 (2024).
11.
P. Sharma, B. Sharma, A. Reza, K. K. Inampudi, R. K. Dhamija, A systematic review of retractions in biomedical research publications: reasons for retractions and their citations in Indian affiliations. Humanities and Social Sciences Communications10 (1), 1–12 (2023).
12.
Y. Sevryugina, R. Jimenez, Analysis of Retracted Manuscripts in Chemistry: Errors vs Misconduct. ACS omega8 (35), 31568–31574 (2023).
13.
H. Peng, D. M. Romero, E.-Á. Horvát, Dynamics of cross-platform attention to retracted papers. Proceedings of the National Academy of Sciences119 (25), e2119086119 (2022).
14.
C. Candal-Pedreira, et al., Retracted papers originating from paper mills: cross sectional study. bmj379 (2022).
15.
M. C. Ferraro, et al., Characteristics of retracted publications related to pain research: a systematic review. Pain pp. 10–1097 (2022).
16.
I. Heibi, S. Peroni, A qualitative and quantitative analysis of open citations to retracted articles: the Wakefield 1998 et al.’s case. Scientometrics126 (10), 8433–8470 (2021).
17.
S. Yang, F. Qi, H. Diao, I. Ajiferuke, Do retraction practices work effectively? Evidence from citations of psychological retracted articles. Journal of Information Science50 (2), 531–545 (2024).
18.
I. Heibi, S. Peroni, A quantitative and qualitative open citation analysis of retracted articles in the humanities. Quantitative Science Studies3 (4), 953–975 (2022).
19.
R. R. Wadhwa, C. Rasendran, Z. B. Popovic, S. E. Nissen, M. Y. Desai, Temporal trends, characteristics, and citations of retracted articles in cardiovascular medicine. JAMA Network Open4 (7), e2118263–e2118263 (2021).
20.
K.-C. Lin, Y.-C. Chen, M.-H. Lin, T.-J. Chen, The trend and ripple effects of retractions in primary health care: A bibliometric analysis. Journal of the Chinese Medical Association87 (10), 927–932 (2024).
21.
D. Fanelli, J. Wong, D. Moher, What difference might retractions make? An estimate of the potential epistemic cost of retractions on meta-analyses. Accountability in Research29 (7), 442–459 (2022).
22.
A. M. Stern, A. Casadevall, R. G. Steen, F. C. Fang, Financial costs and personal consequences of research misconduct resulting in retracted publications. Elife3, e02956 (2014).
23.
E. M. Bucci, On zombie papers. Cell death & disease10 (3), 189 (2019).
24.
G. Cabanac, Chain retraction: how to stop bad science propagating through the literature. Nature632 (8027), 977–979 (2024).
25.
S. E. Swogger, The Interactive Web–Identifying Retracted Publications. Journal of Electronic Resources in Medical Libraries21 (3), 153–161 (2024).
28.
K. Feakins, et al., RetractoBot: a protocol for a randomised controlled trial to assess the impact of notifying authors that they have cited a retracted paper. figshare (2019), journal contribution, doi:10.6084/m9.figshare.6177014.v4, https://doi.org/10.6084/m9.figshare.6177014.v4.
29.
R. Van Noorden, M. Naddaf, Exclusive: the papers that most heavily cite retracted studies. Nature633 (8028), 13–15 (2024).
30.
A. Avenell, M. J. Bolland, G. D. Gamble, A. Grey, A randomized trial alerting authors, with or without coauthors or editors, that research they cited in systematic reviews and guidelines has been retracted. Accountability in Research31 (1), 14–37 (2024).
31.
R. Sinatra, D. Wang, P. Deville, C. Song, A.-L. Barabási, Quantifying the evolution of individual scientific impact. Science354 (6312), aaf5239 (2016).
32.
A. M. Petersen, et al., Reputation and impact in academic careers. Proceedings of the National Academy of Sciences111 (43), 15316–15321 (2014).
33.
D. Van Dijk, O. Manor, L. B. Carey, Publication metrics and success on the academic job market. Current Biology24 (11), R516–R517 (2014).
34.
C. B. Wright, N. L. Vanderford, What faculty hiring committees want. Nature biotechnology35 (9), 885–887 (2017).
35.
R. Kinney, et al., The semantic scholar open data platform. arXiv preprint arXiv:2301.10140 (2023).
36.
J. Zhan, et al., Evaluatology: The science and engineering of evaluation. BenchCouncil Transactions on Benchmarks, Standards and Evaluations4 (1), 100162 (2024).
40.
D. Wang, C. Song, A.-L. Barabási, Quantifying long-term scientific impact. Science342 (6154), 127–132 (2013).
41.
I. Fuyuno, D. Cyranoski, Cash for papers: putting a premium on publication. Nature441 (7095), 792–793 (2006).
42.
Z. Lin, Y. Yin, L. Liu, D. Wang, SciSciNet: A large-scale open data lake for the science of science research. Scientific Data10 (1), 315 (2023).
Acknowledgments
Funding:
Author contributions:
Competing interests:
∗Corresponding author. Email: zhanjianfeng@ict.ac.cn or gaowanling@ict.ac.cn or zhifeiz@ccmu.edu.cn or haoty@m.scnu.edu.cn
This PDF file includes:
Materials and Methods
Figures S1
Tables S1 to S6
Materials and Methods
Data sources and preparation
@@ Acknowledgments
The Semantic Scholar Database (SSD) (?) covers a wide range of publication records, authors, institutions, entities and various interactions between them (e.g., authorships, citations).
@@
The database contains a variety of paper metadata, such as the title, DOI, authors, journal, field, publication date, and citation count etc.
@@
We obtained the set of retracted articles from Retraction Watch (?). This database encompasses 49,721 papers retracted as of 2024, with each record containing metadata including the title, journal, publication date, and retraction date.
@@
The SciSciNet dataset (?), an extension of the Microsoft Academic Graph (MAG) offering comprehensive author, affiliations, journal metadata and linkage information among these, contained 136,726,948 publications.
@@
Computation of Harm Metric.
For each citing paper belonging to the set of citing papers , we compute a harm metric based on its citation patterns relative to a comparator group . Given a citing paper published in venue in year , and associated with a set of fields of study , the comparator set (denoted as ) is constructed by selecting papers that satisfy the following criteria:
@@
where represents the total citations for the comparator set , and is the total citation count for each paper; represents the yearly citations for the comparator set in the -th year after publication, and is the number of citations received by a paper in the -th year after its publication. Both and are derived from the SSD. The total citation count for each paper is obtained from the SSD and the citations for each year after publication are derived from the citation records within the SSD.
Let denote the number of papers in the comparator set . The harm metric for is then computed by comparing its citation metrics to the average citation metrics of its comparator set :
where represents the Harm Metric for total citations for paper ; represents the Harm Metric for yearly citations for paper in year .
@@
where ; for
This represents the computed harm metric for based on its total and yearly citation patterns relative to its comparator group.
Statistics
Following the computation of the harm metric for each paper, as defined in the previous section, we conduct a statistical analysis to investigate the trends and patterns of this metric across different research domains and in aggregate. This analysis is divided into three main parts.
@@
We analyze the overall and field-specific trends of the harm metric over a 10-year window following each paper’s publication. For both the entire corpus and individual fields, we compute the quartile statistics (Q1, Q2, Q3) of the harm metric distribution:
where represents all papers in the corpus, denotes papers in field , is the set of all fields, and is the harm metric of paper . Note that papers with multiple field affiliations contribute to multiple sets.
@@
We further analyze the harm metric distribution across different journal impact factor (IF) ranges. Papers are categorized into five groups based on their venue’s impact factor: . Similar to the field-based analysis, we compute quartile statistics for each IF group:
@@
Citation Timing Analysis of Harm.
@@
We categorize papers based on when they cited (directly or indirectly) retracted papers relative to the retraction date. Let be a retracted paper and be a paper that cites , either directly or indirectly through citation chains. We divide papers into two groups according to whether their citations occurred before or after the retraction of . Similar to the previous analyses, we compute quartile statistics for two groups of papers. For papers citing multiple retracted papers, we use the earliest retraction date as the reference point.
@@
where represents when paper cited (directly or indirectly) the retracted paper , is the retraction date of paper , and is the set of all retracted papers directly or indirectly cited by paper .
@@
Figure S1: The harm of citing retracted papers, with no duplicate statistics across different citation distances.(A) The median harm experienced by papers with an indirect citation distance of 2 varies over time after publication. (B) The median harm experienced by papers with an indirect citation distance of 3 varies over time after publication. (C) The median harm experienced by papers with an indirect citation distance of 4 varies over time after publication. (D) The median harm experienced by papers with an indirect citation distance of 5 varies over time after publication. (E) The median harm experienced by papers with an indirect citation distance of 6 varies over time after publication.
@@
Table S1: The median harm experienced by papers directly citing retracted articles across different fields varies over time after their publication (Part 1).
@@
Table S1: The median harm experienced by papers directly citing retracted articles across different fields varies over time after their publication (Part 2).
@@
Table S2: Whether top 30 universities emphasize recent research outputs in promotion decisions
Table S3: The median harm experienced by papers indirectly citing retracted articles varies over time across different citation distances.
@@
Table S4: After removing duplicate statistics, the median harm experienced by papers indirectly citing retracted articles varies over time across different citation distances.
@@
Table S5: The median harm experienced by papers citing retracted articles, across different citation distances and published in journals with varying impact factors, varies over time after their publication (Part 1).
@@
3
35
1127176
27.62
(-32.6771.40)
27.07
(-31.1766.34)
29.38
(-28.9969.36)
32.59
(-27.3073.05)
36.12
(-24.8375.30)
38.08
(-23.4576.69)
40.41
(-21.2777.92)
43.58
(-18.5078.90)
46.27
(-16.6382.40)
48.62
(-14.95100.00)
3
510
722828
26.70
(-30.5565.21)
26.04
(-28.9363.78)
28.74
(-27.1166.59)
31.57
(-25.0369.22)
34.71
(-23.0970.89)
36.66
(-21.2573.62)
39.27
(-19.2676.55)
42.21
(-17.1578.56)
44.24
(-15.2380.19)
45.99
(-13.5381.04)
Table S5: The median harm experienced by papers citing retracted articles, across different citation distances and published in journals with varying impact factors, varies over time after their publication (Part 2).
@@
6
35
1707950
33.46
(-26.0874.02)
32.58
(-24.4873.19)
35.14
(-22.7075.01)
38.06
(-20.6876.55)
40.29
(-19.2677.94)
42.70
(-17.1378.79)
44.86
(-15.5380.05)
48.17
(-13.4783.16)
50.16
(-11.58100.00)
53.08
(-9.53100.00)
6
510
932141
31.58
(-25.2869.31)
30.84
(-23.6468.98)
33.85
(-21.9970.83)
36.50
(-19.9872.64)
38.85
(-18.0375.48)
41.31
(-16.2778.52)
43.69
(-14.5579.99)
45.86
(-12.4281.40)
47.93
(-10.4982.35)
50.27
(-9.3283.97)
Table S6: Before and after retraction, the median harm experienced by papers citing retracted articles varies over time across different citation distances.

### v2

@@ 3 Discussion
Our findings indicate that the harm caused by retracted papers and their dissemination is closely linked to entrenched “norms” in academia. Specifically, the excessive emphasis on recent research while neglecting long-term contributions, combined with the reliance on journal impact factors as primary evaluation metrics, prevents stakeholders from effectively addressing the harm caused by retracted papers. In the absence of punitive measures, addressing these enduring “norms” within the academic system becomes critically important. Our study connects the harm caused by problematic papers to systemic challenges within the academic recognition framework, offering an additional perspective and direction for addressing this issue.
Our study has certain limitations. The forms of harm are not limited to paper citations; they may also include factors such as the likelihood of publication, funding acquisition, and reputation. However, quantifying aspects such as reputation and funding is challenging, and data availability is limited. Meanwhile, the publication and citation of papers are widely recognized as the “currency” in the academic community and are closely linked to other factors like reputation and funding acquisition [24]. Therefore, analyzing the citation patterns of papers can largely reflect the harm caused by retracted papers. Additionally, due to the lack of data on the paper submission process, we were unable to directly connect publication with the harm caused by retracted papers or explore the potential harm associated with the publication process. However, citation count remains crucial in academic evaluation, effectively revealing both the harm caused by retracted papers and the propagation of that harm. Our findings have broad policy implications, offering valuable insights for policy development. Citation-based evaluation mechanisms are currently widespread across various stakeholders (e.g., institutions, authors, journals), influencing reward programs, research funding allocation, award decisions, and even salary and bonus structures, as well as journal rankings [41].
@@
4 Methods
4.1 Data sources and preparation
@@
The Semantic Scholar Database (SSD) [35] covers a wide range of publication records, authors, institutions, entities and various interactions between them (e.g., authorships, citations).
@@
The database contains a variety of paper metadata, such as the title, DOI, authors, journal, field, publication date, citation count, etc.
@@
We obtained the set of retracted articles from Retraction Watch [7]. This database encompasses 49,721 papers retracted as of 2024, with each record containing metadata including the title, journal, publication date, and retraction date.
@@
The SciSciNet dataset [42], an extension of the Microsoft Academic Graph (MAG) offering comprehensive author, affiliations, journal metadata and linkage information among these, contained 136,726,948 publications.
@@ 4.2 Quantification of harm
Harm Measurement.
For each citing belonging to the set of citing papers , we compute a harm based on its citation patterns relative to a comparator group . Given a citing published in venue in year , and associated with a set of fields of study , the comparator set (denoted as ) is constructed by selecting papers that satisfy the following criteria:
@@
where represents the total citations for the comparator set , is the total citation count for each paper, represents the yearly citations for the comparator set in the -th year after publication, and is the number of citations received by a paper in the -th year after its publication. Both and are derived from the SSD. The total citation count for each paper is obtained from the SSD and the citations for each year after publication are derived from the citation records within the SSD.
Let denote the number of papers in the comparator set . The harm for is then computed by comparing its citation metrics to the average citation metrics of its comparator set :
where represents the harm for total citations for paper , represents the harm for yearly citations for paper in year .
@@
where,
This represents the computed harm for based on its total and yearly citation patterns relative to its comparator group.
4.3 Statistical analysis
Following the computation of the harm for each paper , as defined in the previous section, we conduct a statistical analysis to investigate the trends and patterns of this metric across different research domains and in aggregate. This analysis is divided into three main parts.
@@ 4.3 Statistical analysis
We analyze the overall and field-specific trends of the harm over a 10-year window following each paper’s publication. For both the entire corpus and individual fields, we compute the quartile statistics (Q1, Q2, Q3) of the harm distribution:
where represents all papers in the corpus, denotes papers in field , is the set of all fields, and is the harm of paper . Note that papers with multiple field affiliations contribute to multiple sets.
@@
We further analyze the harm distribution across different journal impact factor (IF) ranges. Papers are categorized into five groups based on their venue’s impact factor: , and . Similar to the field-based analysis, we compute quartile statistics for each IF group:
@@
Harm Analysis of Papers Citing Retracted Papers Before and After Retraction.
@@
We categorize papers based on the date they cited (directly or indirectly) retracted papers relative to the retraction date. Let be a retracted paper and be a paper that cites , either directly or indirectly through citation chains. We divide papers into two groups according to whether their citations occurred before or after the retraction of . Similar to the previous analyses, we compute quartile statistics for two groups of papers. For papers citing multiple retracted papers, we use the earliest retraction date as the reference point.
@@
where represents the date paper cited (directly or indirectly) the retracted paper , is the retraction date of , and is the set of all retracted papers directly or indirectly cited by paper .
Declarations
Funding.
Competing interests.
Author contributions.
Piller [2022]
Piller, C.:
Blots on a field?
Science
377(6604),
358–363
Brainard et al. [2018]
Brainard, J.,
You, J., et al.:
What a massive database of retracted papers reveals about science publishing’s ‘death penalty’.
Science
25(1),
1–5
Else [2024]
Else, H.:
Biomedical paper retractions have quadrupled in 20 years—why?
Nature
630(8016),
280–281
MAURYA and MUNSHI [2024]
MAURYA, M.,
MUNSHI, R.:
Reasons for retraction of clinical research articles in pubmed indexed medical journals from 2012 to 2022.
Med Ethics
9(4),
296–300
Sharma et al. [2023]
Sharma, P.,
Sharma, B.,
Reza, A.,
Inampudi, K.K.,
Dhamija, R.K.:
A systematic review of retractions in biomedical research publications: reasons for retractions and their citations in indian affiliations.
Humanities and Social Sciences Communications
10(1),
1–12
Sevryugina and Jimenez [2023]
Sevryugina, Y.,
Jimenez, R.:
Analysis of retracted manuscripts in chemistry: Errors vs misconduct.
ACS omega
8(35),
31568–31574
Peng et al. [2022]
Peng, H.,
Romero, D.M.,
Horvát, E.-Á.:
Dynamics of cross-platform attention to retracted papers.
Proceedings of the National Academy of Sciences
119(25),
2119086119
Candal-Pedreira et al. [2022]
Candal-Pedreira, C.,
Ross, J.S.,
Ruano-Ravina, A.,
Egilman, D.S.,
Fernández, E.,
Pérez-Ríos, M.:
Retracted papers originating from paper mills: cross sectional study.
bmj
379
Ferraro et al. [2022]
Ferraro, M.C.,
Moore, R.A.,
C Williams, A.C.,
Fisher, E.,
Stewart, G.,
Ferguson, M.C.,
Eccleston, C.,
O’Connell, N.E.:
Characteristics of retracted publications related to pain research: a systematic review.
Pain,
10–1097
Heibi and Peroni [2021]
Heibi, I.,
Peroni, S.:
A qualitative and quantitative analysis of open citations to retracted articles: the wakefield 1998 et al.’s case.
Scientometrics
126(10),
8433–8470
Yang et al. [2024]
Yang, S.,
Qi, F.,
Diao, H.,
Ajiferuke, I.:
Do retraction practices work effectively? evidence from citations of psychological retracted articles.
Journal of Information Science
50(2),
531–545
Heibi and Peroni [2022]
Heibi, I.,
Peroni, S.:
A quantitative and qualitative open citation analysis of retracted articles in the humanities.
Quantitative Science Studies
3(4),
953–975
Wadhwa et al. [2021]
Wadhwa, R.R.,
Rasendran, C.,
Popovic, Z.B.,
Nissen, S.E.,
Desai, M.Y.:
Temporal trends, characteristics, and citations of retracted articles in cardiovascular medicine.
JAMA Network Open
4(7),
2118263–2118263
Lin et al. [2024]
Lin, K.-C.,
Chen, Y.-C.,
Lin, M.-H.,
Chen, T.-J.:
The trend and ripple effects of retractions in primary health care: A bibliometric analysis.
Journal of the Chinese Medical Association
87(10),
927–932
Fanelli et al. [2022]
Fanelli, D.,
Wong, J.,
Moher, D.:
What difference might retractions make? an estimate of the potential epistemic cost of retractions on meta-analyses.
Accountability in Research
29(7),
442–459
Stern et al. [2014]
Stern, A.M.,
Casadevall, A.,
Steen, R.G.,
Fang, F.C.:
Financial costs and personal consequences of research misconduct resulting in retracted publications.
Elife
3,
02956
Bucci [2019]
Bucci, E.M.:
On zombie papers.
Cell death & disease
10(3),
189
Cabanac [2024]
Cabanac, G.:
Chain retraction: how to stop bad science propagating through the literature.
Nature
632(8027),
977–979
Swogger [2024]
Swogger, S.E.:
The interactive web–identifying retracted publications.
Journal of Electronic Resources in Medical Libraries
21(3),
153–161
Feakins et al. [2019]
Feakins, K.,
Irving, F.,
DeVito, N.,
Feakins, B.,
Curtis, H.,
Bacon, S., et al.:
RetractoBot: a protocol for a randomised controlled trial to assess the impact of notifying authors that they have cited a retracted paper.
figshare
https://doi.org/10.6084/m9.figshare.6177014.v4 .
Journal contribution
Van Noorden and Naddaf [2024]
Van Noorden, R.,
Naddaf, M.:
Exclusive: the papers that most heavily cite retracted studies.
Nature
633(8028),
13–15
Avenell et al. [2024]
Avenell, A.,
Bolland, M.J.,
Gamble, G.D.,
Grey, A.:
A randomized trial alerting authors, with or without coauthors or editors, that research they cited in systematic reviews and guidelines has been retracted.
Accountability in Research
31(1),
14–37
Sinatra et al. [2016]
Sinatra, R.,
Wang, D.,
Deville, P.,
Song, C.,
Barabási, A.-L.:
Quantifying the evolution of individual scientific impact.
Science
354(6312),
5239
Petersen et al. [2014]
Petersen, A.M.,
Fortunato, S.,
Pan, R.K.,
Kaski, K.,
Penner, O.,
Rungi, A.,
Riccaboni, M.,
Stanley, H.E.,
Pammolli, F.:
Reputation and impact in academic careers.
Proceedings of the National Academy of Sciences
111(43),
15316–15321
Van Dijk et al. [2014]
Van Dijk, D.,
Manor, O.,
Carey, L.B.:
Publication metrics and success on the academic job market.
Current Biology
24(11),
516–517
Wright and Vanderford [2017]
Wright, C.B.,
Vanderford, N.L.:
What faculty hiring committees want.
Nature biotechnology
35(9),
885–887
Kinney et al. [2023]
Kinney, R.,
Anastasiades, C.,
Authur, R.,
Beltagy, I.,
Bragg, J.,
Buraczynski, A.,
Cachola, I.,
Candra, S.,
Chandrasekhar, Y.,
Cohan, A., et al.:
The semantic scholar open data platform.
arXiv preprint arXiv:2301.10140
Zhan et al. [2024]
Zhan, J., et al.:
Evaluatology: The science and engineering of evaluation.
BenchCouncil Transactions on Benchmarks, Standards and Evaluations
4(1),
100162
Samoylenko et al. [2023]
Samoylenko, I.,
Aleja, D.,
Primo, E.,
Alfaro-Bittner, K.,
Vasilyeva, E.,
Kovalenko, K.,
Musatov, D.,
Raigorodskii, A.M.,
Criado, R.,
Romance, M.,
Papo, D.,
Perc, M.,
Barzel, B.,
Boccaletti, S.:
Why are there six degrees of separation in a social network?
Phys. Rev. X
13,
021032
https://doi.org/10.1103/PhysRevX.13.021032
Wang et al. [2013]
Wang, D.,
Song, C.,
Barabási, A.-L.:
Quantifying long-term scientific impact.
Science
342(6154),
127–132
Fuyuno and Cyranoski [2006]
Fuyuno, I.,
Cyranoski, D.:
Cash for papers: putting a premium on publication.
Nature
441(7095),
792–793
Lin et al. [2023]
Lin, Z.,
Yin, Y.,
Liu, L.,
Wang, D.:
Sciscinet: A large-scale open data lake for the science of science research.
Scientific Data
10(1),
315
@@
Extended Data Fig. 1: The harm of citing retracted papers, with no duplicate statistics across different citation distances.(a) The median harm experienced by papers with an indirect citation distance of 2 varies over time after publication. (b) The median harm experienced by papers with an indirect citation distance of 3 varies over time after publication. (c) The median harm experienced by papers with an indirect citation distance of 4 varies over time after publication. (d) The median harm experienced by papers with an indirect citation distance of 5 varies over time after publication. (e) The median harm experienced by papers with an indirect citation distance of 6 varies over time after publication.
@@
Extended Data Table 1: The median harm experienced by papers directly citing retracted articles across different fields varies over time after their publication (Part 1).
@@
Geography
38
45.40
(-47.91100.00)
41.22
(-30.18100.00)
48.69
75.00
(-3.38100.00)
72.41
Y6 Harm (IQR)
Y7 Harm (IQR)
Y8 Harm (IQR)
Y9 Harm (IQR)
Y10 Harm (IQR)
26.49
(-55.1375.61)
30.43
(-51.7578.46)
33.44
(-48.2381.69)
37.21
(-44.1089.09)
40.08
(-41.39100.00)
20.59
(-60.1772.26)
26.09
(-48.4875.91)
27.44
(-53.1277.33)
31.66
(-43.1080.24)
32.33
(-43.6986.65)
14.29
(-72.04100.00)
46.19
(-94.59100.00)
40.59
(-200.00100.00)
56.64
(-96.63100.00)
100.00
(-18.45100.00)
23.78
(-52.7369.45)
27.79
(-50.5172.48)
30.97
(-46.5975.12)
33.71
(-42.9077.07)
36.38
(-40.6280.80)
66.67
(-25.00100.00)
60.49
(-39.83100.00)
62.24
(-31.96100.00)
82.12
(-18.90100.00)
100.00
(-12.35100.00)
27.82
(-47.7772.47)
32.80
(-46.5175.07)
34.44
(-44.2577.73)
37.14
(-40.5881.01)
39.13
(-37.2588.08)
75.28
(-12.10100.00)
78.48
(-17.30100.00)
84.16
(-7.81100.00)
100.00
100.00
58.20
(-25.69100.00)
62.08
(-13.48100.00)
78.45
91.78
84.21
44.00
(-59.78100.00)
48.91
(-36.09100.00)
55.74
(-42.20100.00)
59.38
(-29.31100.00)
74.85
(-22.11100.00)
42.82
(-38.65100.00)
45.47
(-41.41100.00)
53.32
(-31.75100.00)
58.66
(-22.07100.00)
65.46
(-18.40100.00)
@@
Extended Data Table 1: The median harm experienced by papers directly citing retracted articles across different fields varies over time after their publication (Part 2).
@@
Extended Data Table 2: Whether top 30 universities emphasize recent research outputs in promotion decisions
Extended Data Table 3: The median harm experienced by papers indirectly citing retracted articles varies over time across different citation distances.
@@
Extended Data Table 4: After removing duplicate statistics, the median harm experienced by papers indirectly citing retracted articles varies over time across different citation distances.
@@
Extended Data Table 5: The median harm experienced by papers citing retracted articles, across different citation distances and published in journals with varying impact factors, varies over time after their publication (Part 1).
@@
Extended Data Table 5: The median harm experienced by papers citing retracted articles, across different citation distances and published in journals with varying impact factors, varies over time after their publication (Part 2).
@@
Extended Data Table 6: Before and after retraction, the median harm experienced by papers citing retracted articles varies over time across different citation distances.